OpenStudy (anonymous):

Of the plants in Mr. Jacob's garden, 30% are tomatoes, 10% are radishes, 15% are cucumbers, and the remaining 9 are carrots. What is the total # of plants in his garden? 9 20 36 50 59

OpenStudy (johnweldon1993):

Well lets see, we know that it all adds up to 100% \[\large 30 + 10 + 15 = 55\] So far we have 55 percent of the garden...the remaining 9 plants are carrots If we have 55% NOT including carrots, that means that the remaining 45% ARE carrots, so right away we know that 45% = 9 since the whole carrot crop is only 9 plants Alright, so lets break it down 30% = tomatoes 15% = cucumbers 10% = radishes 45% = carrots = 9 So if 45% = 9 How many times does 15% go into 45%? well 3 times right? So that means that there would be 3 times less cucumbers than carrots... 9 / 3 = 3 So there are 3 cucumbers. 30% = tomatoes 15% = cucumbers = 3 10% = radishes 45% = carrots = 9 Now how many times does 15% go into 30%? Well 2 times right? That means there are 2 times MORE tomatoes than there are cucumbers. 3 * 2 = 6 So there are 6 tomatoes 30% = tomatoes = 6 15% = cucumbers = 3 10% = radishes 45% = carrots = 9 And finally how many times does 10 go into 30? well 3 right? So that means there are 3 times LESS radishes than tomatoes 6 / 3 = 2 So there are 2 radishes 30% = tomatoes = 6 15% = cucumbers = 3 10% = radishes = 2 45% = carrots = 9 Now altogether, we have 6 + 3 + 2 + 9 = 20 plants in the garden

OpenStudy (triciaal):

since they did not ask for the breakdown it is a waste of time to do all that. additionally, you don't want to round. now as at the start, 55% not carrots therefore 45% of the total, x = 9 45/100* x = 9 x = 900/45 = 20